# Graph Masked Autoencoder for Spatio-Temporal Graph Learning

### v1

@@ III-D The Layout of Spatial-Temporal Graph Mask Autoencoder (STGMAE)
Given the spatio-temporal graph $\mathcal{G}$, the encoder function $f_{E}$ and the decoder function $f_{D}$, Graph mask autoencoder (GMAE) aims to reconstruct the spatio-temporal graph $\mathcal{G}^{\prime}$ as follows:
$\textbf{H}=f_{E}(\tilde{\mathbf{A}},\tilde{\mathbf{X}}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{G}^{\prime}=f_{D}(\tilde{\mathbf{A}}^{\prime},\tilde{\mathbf{H}})$ (4)
@@
where $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{X}}$ denote the masked adjacent matrix and mask node embedding matrix respectively. And $\tilde{\mathbf{A}^{\prime}}$ and $\tilde{\mathbf{H}}$ represent the remasked adjacent matrix and the remasked node embedding matrix respectively. Details of GMAE design for the spatio-temporal graph $\mathcal{G}$ are in the following sections.
Input: The heterogeneous spatial-temporal graph $\mathcal{G}$, the maximum epoch number $E$, the learning rate $\eta$;
Output: Regional Embedding $\mathbf{H}$
and trained parameters in $\mathbf{\Theta}$;
1 Initialize all parameters in $\mathbf{\Theta}$;
2 Train the framework STGMAE by Equation 9
3 for $epoch=1,2,...,E$ do
4       Random select some nodes and mask them according to Equation 6;
5       Random select some links and mask them;
6       Send the masked nodes features and the adjacent matrix to GCN Encoder;
7       Remark the node features and the adjacent matrix and send them to GCN Decoder;
8       Compute the cosine similarity loss $\mathcal{L}_{1}$ following Equation 8 and MSE loss $\mathcal{L}_{2}$ according to Equation 8;
9       Minimize the loss $\mathcal{L}_{rec}$ by Equation 9 using gradient decent with learning rate $\eta$;
10       for $\theta\in\mathbf{\Theta}$ do
11             $\theta=\theta-\eta\cdot\frac{\partial\mathcal{L}_{rec}}{\partial\theta}$
12       end for
13
14 end for
Return $\mathbf{H}$ and all parameters $\mathbf{\Theta}$;
Algorithm 1 The STGMAE Learning Algorithm
III-D1 Motivation of Features and Structure Reconstruction
@@
$\begin{matrix}\underbrace{\mathcal{L}}\\ \text{Graph}\end{matrix}\quad=\quad\begin{matrix}\underbrace{\|\mathbf{X}-\hat{\mathbf{X}}\|_{F}^{2}}\\ \text{Node}\end{matrix}+\quad\alpha\cdot\begin{matrix}\underbrace{\|\mathbf{A}-\sigma(\mathbf{H}\cdot\mathbf{H^{T}})}\|_{F}^{2}\\ \text{Structure}\end{matrix}$ (5)
However, GAE often prioritizes reconstructing the features $\mathbf{X}$ while minimizing or ignoring the reconstruction error related to the structure $\mathbf{A}$. Actually, by adapting $\alpha$ based on the specific task and application, GAEs can effectively learn representations that cater to the demands of diverse graph-based problems. Inspired by recent studies [20, 31] and the significance of different types in our heterogeneous graph, we propose to reconstruct both the structure and features in our STGMAE. Besides, we use a mask mechanism-based generative self-supervised learning paradigm, which allows for effective learning of the spatio-temporal dependencies of region-wise within the graph.
@@
III-D2 Reconstructing Features and Structures using Mask Mechanism
@@
$x_{i}=\begin{cases}x_{M},&\text{if }v_{i}\in\widetilde{\mathcal{V}}\\ x_{i},&\text{if }v_{i}\notin\widetilde{\mathcal{V}}\end{cases}$ (6)
Given the masked node feature matrix $\widetilde{\mathbf{X}}$, one of the objectives of STGMAE is to reconstruct masked node features of regions (e.g., $r_{i}$) in $\widetilde{\mathcal{V}}$ and masked the adjacent matrix $\widetilde{\mathbf{A}}$ via a learned vector $a_{M}$. In GNNs, the hidden embedding vector of each node is affected by its neighboring nodes’ embedding vectors according to the message passing mechanism. However, selecting nodes via scores can lead to bias, for example, if all neighbors of one node are selected or masked. To prevent bias in node selection, random node selection or sampling is taken into account the entire graph structure rather than just local node properties, which also improves robustness.
III-D3 Remask Decoder via GNNs
The role of the decoder function $f_{D}$ is to map the hidden embedding vectors $\tilde{\mathbf{H}}$ to the input graph $\mathcal{G}^{\prime}$ and its node embedding matrix $\mathbf{X}^{\prime}$, based on the semantic level of the target input graph $\mathcal{G}$ and its corresponding node embedding matrix $\mathbf{X}$ [32]. For example, in NLP, an MLP is typically sufficient for targets of one-hot missing words with rich semantics, as there is no structure semantics to consider. However, in CV, a transformer is often used as the decoder to recover pixel spots with few semantics.
@@ III-D3 Remask Decoder via GNNs
To further improve the ability of the encoder to learn compressed graph representations, we adopt a remask mechanism in the decoder, inspired by [21]. The process is represented as following:
$\widetilde{h}_{i}=\begin{cases}h_{M}&\leavevmode\nobreak\ \leavevmode\nobreak\ v_{i}\in\widetilde{\mathcal{V}}\\ h_{i}&v_{i}\notin\widetilde{\mathcal{V}}\end{cases}$ (7)
Here, $\widetilde{\mathbf{H}}=REMASK(\mathbf{H})$, where a masked node is forced to build its input feature based on the neighboring non-masked nodes’ hidden embedding vectors, under the guidance of the GCN decoder. This approach helps ensure that the learned embeddings are both informative and compressed, which can improve the overall region representation ability of our framework STGMAE.
III-E Model Optimization and Complexity
To optimize our method STGMAE, we adopt the cosine similarity loss, denoted as $\mathcal{L}_{1}$, to optimize remasked node features and initial input node features. Besides, to help optimize structure masking, another loss MSE is adopted. These two losses are shown as:
$\mathcal{L}_{1}$ $=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{h_{i}^{T}\,\widetilde{h}_{i}}{\|\widetilde{h}_{i}\|\cdot\|h_{i}\|}\right);$ (8)
@@ III-E Model Optimization and Complexity
By leveraging generative self-supervised learning tasks during pretraining, our proposed framework, STGMAE, aims to produce high-quality region embeddings. The optimization goal of our framework STGMAE can be formalized as follows:
@@
$\min\limits_{\theta}\mathcal{L}_{rec}(\mathbf{A},\mathbf{X},\widetilde{\mathbf{A}}^{\prime},\widetilde{\mathbf{X}}^{\prime}),\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{L}_{rec}=\mathcal{L}_{1}+\mathcal{L}_{2}$ (9)
@@
where $\theta$ denotes the parameters of the model, $\mathcal{L}_{rec}$ is the reconstruction loss that measures the difference between the original and reconstructed adjacency matrix $\mathbf{A}$ and node feature matrix $\mathbf{X}$. The objective function ensures that the model can accurately reconstruct both the graph structure and node features, while the second term encourages the model to learn representations that are useful for predicting node features in a self-supervised manner. Overall, the optimization goal of our framework reflects our aim to learn high-quality region embeddings that can be used for downstream tasks such as node classification, link prediction, and community detection. And the time complexity of our framework STGMAE is divided into two parts: first of all, the multi-view space-time message passing paradigm has a complexity of $\mathcal{O}(|\mathcal{V}|\times L\times d)$, where $N$ and $L$ denote the number of edges and graph propagation layers, respectively. Secondly, for loss calculation, STGMAE takes $\mathcal{O}(S\times|\tilde{\mathcal{V}}|\times d)$, where $S$ denotes the number of sampled nodes included in each batch for the spatio-temporal graph.
III-F In-depth Theory Analysis of STGMAE
In this section, we aim to provide in-depth analysis of our framework STGMAE. We present a small STGMAE loss denoting a small spatio-temporal alignment loss. Firstly, we have the following assumption: There exists a pseudo-inverse encoder $f_{E}$ which also the resulting pseudo autoencoder $s_{g}=\left\{g,f_{g}\right\}$satisfies $\mathbf{E}||s_{g}(z)-z||\leq\epsilon$, where $z$ denotes masked representation of the input graph $\mathcal{G}$. Following this assumption, our method performs input-output alignment on the masked spatio-temporal graph. The lower bound of $\mathcal{L}_{rec}\geq\mathcal{L}_{asym}-\epsilon+C$, where $C$ is a constant number and $\mathcal{L}_{asym}$ is presented as following for the spatio-temporal graph $\mathcal{G}$:
$\mathcal{L}_{asym}=-\mathbf{E}_{x_{1},x_{2}}f(x_{1})s_{g}(x_{2})=-tr(\mathbf{H}_{g}^{T}\bar{\mathbf{A}}_{m}\mathbf{H})$ (10)
@@ III-F In-depth Theory Analysis of STGMAE
where $\mathbf{H}$ is the output matrix of our method STGMAE on $\mathcal{G}_{1}$ whose $x_{1}$-row is $\mathbf{H}_{x_{1}}=\sqrt{d_{x_{1}}}f(x_{1})$, and $\mathbf{H}_{g}$ is the output matrix of $s_{g}$ on $\mathcal{G}_{2}$ whose $x_{2}$-th row is $(\mathbf{H}_{g})_{x_{2}}=\sqrt{d_{x_{2}}}s_{g}(x_{2})$. Thus, a small STGMAE loss implies a small alignment loss as a lower bound, which also verifies align mask and unmask view via STGMAE. This again verifies that our method can obtain the same level performance as the contrastive learning loss without data augmentation.
Table I: Data Description of Experimented Datasets
@@
IV Evaluation
We conduct experiments to answer the following questions: RQ1: How does the STGMAE fare in terms of different baselines across multiple spatial-temporal learning applications including traffic prediction, crime prediction and house price prediction? RQ2: How do various data perspectives and generative elements impact the effectiveness of region representation? RQ3: How well does our STGMAE fare in representation learning over areas with various levels of data sparsity? RQ4: What effects do different hyperparameter settings have on the performance of the STGMAE’s region representation? RQ5: How the efficiency our method STGMAE than other region representation methods (e.g., MV-PN, MVURE and MGFN)? RQ6: What advantages do our STGMAE using learned representations across regions have?
IV-A Experimental Setup
IV-A1 Datasets and Protocols
With the aim of assessing the effectiveness of our STGMAE framework, we gathered real-world datasets from two major cities, namely Chicago and New York City. These datasets were utilized to perform three spatio-temporal mining tasks: crime prediction and traffic flow forecasting, as well as property price prediction. To ensure a comprehensive evaluation, we included various types of crimes in our datasets. For the Chicago database, crimes such as Theft, Battery, Assault, and Damage were considered, while the NYC database encompassed crimes such as Burglary, Larceny, Robbery, and Assault. The selection of crime types aligns with the parameters in [33]. A detailed description of the data is shown in Table I.
IV-A2 Hyperparameter Settings
Following existing region representation studies [27, 24], we set the dimensionality $d$ as 96. And from the hyperparameter study experiments, we find that the framework STGMAE obtains the best performance when the GCN depth is set as 2. And the learning rate is 0.001 with weight decay 0.0005. For crime prediction backbone model, the learning rate of ST-SHN is set as 0.001 and weight decay is 0.96. The depth of the spatial aggregation layer is set as 2. For traffic prediction backbone model ST-GCN, the input length is 12 time intervals with each interval setting as 5 minutes. And the output is 15 minutes’ results of traffic flow prediction. Most of baselines are utilized following their released codes.
IV-A3 Implementation Details
To be fair, we implement STGMAE and other baselines in Python 3.8, pytorch 1.7.0 with cuda 11.3 for our method implementation and tensorflow 1.15.3 for traffic prediction and crime prediction. The experiments are conducted on a server with 10-cores of Intel(R) Core(TM) i9-9820X CPU @ 3.30GHz 64.0GB RAM and one Nvidia GeForce RTX 3090 GPU.
Refer to caption Refer to caption
Figure 2: Ablation study of STGMAE on traffic prediction
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: Ablation study of STGMAE on crime prediction
Table II: Overall performance in crime forecasting, traffic prediction, and house price prediction.

### v2

@@ 2.4. The Layout of Spatial-Temporal Graph Mask Autoencoder
Given the spatial-temporal graph $\mathcal{G}$, the encoder function $f_{E}$ and the decoder function $f_{D}$, Graph mask autoencoder (GMAE) aims to reconstruct the spatial-temporal graph $\mathcal{G}^{\prime}$ as follows:
[…]
@@
where $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{X}}$ denote the masked adjacent matrix and mask node embedding matrix respectively. And $\tilde{\mathbf{A}^{\prime}}$ and $\tilde{\mathbf{H}}$ represent the remasked adjacent matrix and the remasked node embedding matrix respectively. Details of GMAE design for the spatial-temporal graph $\mathcal{G}$ are in the following sections.
2.4.1. Motivation of Features and Structure Reconstruction
@@
(5) $\begin{matrix}\underbrace{\mathcal{L}}\\ \text{Graph}\end{matrix}\quad=\quad\begin{matrix}\underbrace{\|\mathbf{X}-\hat{\mathbf{X}}\|_{F}^{2}}\\ \text{Node}\end{matrix}+\quad\alpha\cdot\begin{matrix}\underbrace{\|\mathbf{A}-\sigma(\mathbf{H}\cdot\mathbf{H^{T}})}\|_{F}^{2}\\ \text{Structure}\end{matrix}$
However, GAE often prioritizes reconstructing the features $\mathbf{X}$ while minimizing or ignoring the reconstruction error related to the structure $\mathbf{A}$. Actually, by adapting $\alpha$ based on the specific task and application, GAEs can effectively learn representations that cater to the demands of diverse graph-based problems. Inspired by recent studies (Park et al., 2019; Salehi and Davulcu, 2019) and the significance of different types in our heterogeneous graph, we propose to reconstruct both the structure and features in our HGAurban. Besides, we use a mask mechanism-based generative self-supervised learning paradigm, which allows for effective learning of the spatial-temporal dependencies of region-wise within the graph.
@@
2.4.2. Reconstructing Features and Structures using Mask Mechanism
@@
(6) $x_{i}=\begin{cases}x_{M},&\text{if }v_{i}\in\widetilde{\mathcal{V}}\\ x_{i},&\text{if }v_{i}\notin\widetilde{\mathcal{V}}\end{cases}$
Given the masked node feature matrix $\widetilde{\mathbf{X}}$, one of the objectives of HGAurban is to reconstruct masked node features of regions (e.g., $r_{i}$) in $\widetilde{\mathcal{V}}$ and masked the adjacent matrix $\widetilde{\mathbf{A}}$ via a learned vector $a_{M}$. In GNNs, the hidden embedding vector of each node is affected by its neighboring nodes’ embedding vectors according to the message passing mechanism. However, selecting nodes via scores can lead to bias, for example, if all neighbors of one node are selected or masked. To prevent bias in node selection, random node selection or sampling is taken into account the entire graph structure rather than just local node properties, which also improves robustness.
2.4.3. Remask Decoder via GNNs
The role of the decoder function $f_{D}$ is to map the hidden embedding vectors $\tilde{\mathbf{H}}$ to the input graph $\mathcal{G}^{\prime}$ and its node embedding matrix $\mathbf{X}^{\prime}$, based on the semantic level of the target input graph $\mathcal{G}$ and its corresponding node embedding matrix $\mathbf{X}$ (He et al., 2022). For example, in NLP, an MLP is typically sufficient for targets of one-hot missing words with rich semantics, as there is no structure semantics to consider. However, in CV, a transformer is often used as the decoder to recover pixel spots with few semantics.
@@ 2.4.3. Remask Decoder via GNNs
To further improve the ability of the encoder to learn compressed graph representations, we adopt a remask mechanism in the decoder, inspired by (Hou et al., 2022). The process is represented as following:
(7) $\widetilde{h}_{i}=\begin{cases}h_{M}&\leavevmode\nobreak\ \leavevmode\nobreak\ v_{i}\in\widetilde{\mathcal{V}}\\ h_{i}&v_{i}\notin\widetilde{\mathcal{V}}\end{cases}$
Here, $\widetilde{\mathbf{H}}=REMASK(\mathbf{H})$, where a masked node is forced to build its input feature based on the neighboring non-masked nodes’ hidden embedding vectors, under the guidance of the GCN decoder. This approach helps ensure that the learned embeddings are both informative and compressed, which can improve the overall region representation ability of our framework HGAurban.
2.5. Model Optimization and Complexity
To optimize our method HGAurban, we adopt the cosine similarity loss, denoted as $\mathcal{L}_{1}$, to optimize remasked node features and initial input node features. Besides, to help optimize structure masking, another loss MSE is adopted. These two losses are shown as:
(8) $\mathcal{L}_{1}$ $=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{h_{i}^{T}\,\widetilde{h}_{i}}{\|\widetilde{h}_{i}\|\cdot\|h_{i}\|}\right);$
@@ 2.5. Model Optimization and Complexity
By leveraging generative self-supervised learning tasks during pretraining, our proposed framework, HGAurban, aims to produce high-quality region embeddings. The optimization goal of our framework HGAurban can be formalized as follows:
@@
(9) $\min\limits_{\theta}\mathcal{L}_{rec}(\mathbf{A},\mathbf{X},\widetilde{\mathbf{A}}^{\prime},\widetilde{\mathbf{X}}^{\prime}),\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{L}_{rec}=\mathcal{L}_{1}+\mathcal{L}_{2}$
@@
where $\theta$ denotes the parameters of the model, $\mathcal{L}_{rec}$ is the reconstruction loss that measures the difference between the original and reconstructed adjacency matrix $\mathbf{A}$ and node feature matrix $\mathbf{X}$. The objective function ensures that the model can accurately reconstruct both the graph structure and node features, while the second term encourages the model to learn representations that are useful for predicting node features in a self-supervised manner. Overall, the optimization goal of our framework reflects our aim to learn high-quality region embeddings that can be used for downstream tasks such as node classification, link prediction, and community detection. And the time complexity of our framework HGAurban is divided into two parts: first of all, the multi-view space-time message passing paradigm has a complexity of $\mathcal{O}(|\mathcal{V}|\times L\times d)$, where $N$ and $L$ denote the number of edges and graph propagation layers, respectively. Secondly, for loss calculation, HGAurban takes $\mathcal{O}(S\times|\tilde{\mathcal{V}}|\times d)$, where $S$ denotes the number of sampled nodes included in each batch for the spatial-temporal graph.
2.6. In-depth Theory Analysis of HGAurban
[…]
(10) $\mathcal{L}_{asym}=-\mathbf{E}_{x_{1},x_{2}}f(x_{1})s_{g}(x_{2})=-tr(\mathbf{H}_{g}^{T}\bar{\mathbf{A}}_{m}\mathbf{H})$
@@ 2.6. In-depth Theory Analysis of HGAurban
where $\mathbf{H}$ is the output matrix of our method HGAurban on $\mathcal{G}_{1}$ whose $x_{1}$-row is $\mathbf{H}_{x_{1}}=\sqrt{d_{x_{1}}}f(x_{1})$, and $\mathbf{H}_{g}$ is the output matrix of $s_{g}$ on $\mathcal{G}_{2}$ whose $x_{2}$-th row is $(\mathbf{H}_{g})_{x_{2}}=\sqrt{d_{x_{2}}}s_{g}(x_{2})$. Thus, a small HGAurban loss implies a small alignment loss as a lower bound, which also verifies align mask and unmask view via HGAurban. This again verifies that our method can obtain the same level performance as the contrastive learning loss without data augmentation.
0:
1: Spatio-temporal graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with:
2:  POI matrix $\mathcal{P}\in\mathbb{R}^{I\times C}$
3:  Mobility matrix $\mathbf{M}\in\mathbb{R}^{I\times I\times T}$
4:  Distance matrix $\mathcal{D}\in\mathbb{R}^{I\times I}$
5: Mask ratio $p_{\text{mask}}\in(0,1)$
16: Relation set $\mathcal{R}=\{\mathcal{R}_{p},\mathcal{R}_{m},\mathcal{R}_{d},\mathcal{R}_{pm},\mathcal{R}_{md}\}$
26: Reconstructed adjacency $\hat{\mathbf{A}}$ and features $\hat{\mathbf{X}}$
7:/* Graph Partitioning */
8: Divide urban area into II regions $\{r_{i}\}_{i=1}^{I}$
39: Initialize node features $\mathbf{X}\leftarrow\text{Concat}(\mathbf{P},\mathbf{M},\mathcal{D})$
10:/* Masked Autoencoder Setup */
11: Randomly select $\mathcal{V}_{\text{mask}}\subset\mathcal{V}$ with $|\mathcal{V}_{\text{mask}}|=p_{\text{mask}}|\mathcal{V}|$
412: Generate masked graph $\tilde{\mathcal{G}}$ via Eq. 6 and Eq. 7
13:/* Encoder: Heterogeneous GNN */
14:for $l=1$ to $L$ do
15:  Multi-relational message passing:
16:  $\mathbf{h}_{i}^{(l+1)}\leftarrow\sigma\Big{(}\sum_{\gamma\in\mathcal{R}}\sum_{j\in\mathcal{N}^{\gamma}_{i}}\frac{1}{|\mathcal{N}^{\gamma}_{i}|}\mathbf{W}_{\gamma}^{(l)}\mathbf{h}_{j}^{(l)}\Big{)}$
17:end for
[…]
19:/* Decoder: Reconstruction */
20: Remask embeddings $\tilde{\mathbf{H}}\leftarrow\text{REMASK}(\mathbf{H})$ (Eq. 7)
[…]
23:/* Loss Computation */
24: Feature loss: $\mathcal{L}_{1}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\left(1-\cos(\mathbf{h}_{i},\tilde{\mathbf{h}}_{i})\right)$
25: Structure loss: $\mathcal{L}_{2}\leftarrow\|\mathbf{A}-\sigma(\tilde{\mathbf{A}}^{\prime}\tilde{\mathbf{A}}^{\prime\top})\|_{F}^{2}$
726: Total loss: $\mathcal{L}\leftarrow\mathcal{L}_{1}+\mathcal{L}_{2}$
27:return  $\hat{\mathbf{A}},\hat{\mathbf{X}}$
Algorithm 1 HGAurban: Heterogeneous Graph Autoencoding for Urban Spatial-Temporal Learning
Table 1. Data Description of Experimented Datasets
@@
3. Experiments
We conduct experiments to answer the following questions: RQ1: How does the HGAurban fare in terms of different baselines across multiple spatial-temporal learning applications including traffic prediction, crime prediction and house price prediction? RQ2: How do various data perspectives and generative elements impact the effectiveness of region representation? RQ3: How well does our HGAurban fare in representation learning over areas with various levels of data sparsity? RQ4: What effects do different hyperparameter settings have on the performance of the HGAurban’s region representation? RQ5: How the efficiency our method HGAurban than other region representation methods?
3.1. Experimental Setup
3.1.1. Datasets and Protocols
With the aim of assessing the effectiveness of our HGAurban framework, we gathered real-world datasets from two major cities, namely Chicago and New York City. These datasets were utilized to perform three spatio-temporal mining tasks: crime prediction and traffic flow forecasting, as well as property price prediction. To ensure a comprehensive evaluation, we included various types of crimes in our datasets. For the Chicago database, crimes such as Theft, Battery, Assault, and Damage were considered, while the NYC database encompassed crimes such as Burglary, Larceny, Robbery, and Assault. The selection of crime types aligns with the parameters in (Xia et al., 2021). A detailed description of the data is shown in Table 1.
3.1.2. Hyperparameter Settings
Following existing region representation studies (Wu et al., 2022; Zhang et al., 2021), we set the dimensionality $d$ as 96. And from the hyperparameter study experiments, we find that the framework HGAurban obtains the best performance when the GCN depth is set as 2. And the learning rate is 0.001 with weight decay 0.0005. For crime prediction backbone model, the learning rate of ST-SHN is set as 0.001 and weight decay is 0.96. The depth of the spatial aggregation layer is set as 2. For traffic prediction backbone model ST-GCN, the input length is 12 time intervals with each interval setting as 5 minutes. And the output is 15 minutes’ results of traffic flow prediction.
3.1.3. Implementation Details
To be fair, we implement HGAurban and other baselines in Python 3.8, pytorch 1.7.0 with cuda 11.3 for our method implementation and tensorflow 1.15.3 for traffic prediction and crime prediction. The experiments are conducted on a server with 10-cores of Intel(R) Core(TM) i9-9820X CPU @ 3.30GHz 64.0GB RAM and one Nvidia GeForce RTX 3090 GPU.
Table 2. Overall performance in crime forecasting, traffic prediction, and house price prediction.